OpenStudy (anonymous):

How many years older will you be 1.20 billion seconds from now?

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

How many seconds in a minute?

OpenStudy (anonymous):

60 sec

OpenStudy (skullpatrol):

how many minutes in an hour?

OpenStudy (anonymous):

60 min

OpenStudy (skullpatrol):

how many hours in a day?

OpenStudy (anonymous):

24 hrs ,:p

OpenStudy (skullpatrol):

How many days in a year?

OpenStudy (anonymous):

365:p

OpenStudy (skullpatrol):

1.20 billion seconds = ? years

OpenStudy (anonymous):

hihihi I don't know. ... And the leap-years???? 1 year = 365 days = 24 * 365 hours = 60 * 24 * 365 minuts = 60 * 60 * 24 * 365 seconds = 31536000 years/seconds (1.20 billons /31536000 )seconds = 3.17 x 10^-20 years

OpenStudy (anonymous):

no leap yrs

OpenStudy (skullpatrol):